How many electrons have removed from a positively charged particles if it has a net charge of 5X10-9C?


1
Expert's answer
2022-02-23T08:31:40-0500

The number of electrons was removed from a positively charged particle

"N=q\/e=5*10^{-9}\/1.6*10^{-19}=3.1*10^{10}"